OpenStudy (anonymous):

Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? −b b2 − 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (anonymous):

b^2-4ac does. if it's a perfect square then it's possible to solve by factoring :- 7^2 - (4)(2)(3) = 49 - 24 = 25 25 is a perfect square